OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table?

OpenStudy (anonymous):

P w I 14 2 ? 22 8 ?

OpenStudy (anonymous):

srry table got discombobulated when I pasted xD

OpenStudy (anonymous):

try dividing both sides of equation by 2 and then subtracting w from both sides

OpenStudy (anonymous):

mmmm...k

OpenStudy (anonymous):

just plug in, whr r u stuck?

OpenStudy (anonymous):

cause I'm mathematically challenged thats why xD

OpenStudy (anonymous):

ahh lol

OpenStudy (anonymous):

did u get an expression for l?

OpenStudy (anonymous):

hold on...........

OpenStudy (anonymous):

u just replace the values in formula so u have 14 = 2 * (2+l) 2+l=7 l=5 and 22= 2* (8+l) 8+l=11 l=3

OpenStudy (anonymous):

hey thanks

OpenStudy (anonymous):

replacing the values only completes the second part of the question. The first part has to be solved by dividing by 2 and subtracting w

OpenStudy (anonymous):

you should get l = p/2 - w . This completes first part of the question. Now just substitute the values of p and w and get the answers for l

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

and thanks gps